40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 18 18:13:55 2019, Koji, Update, safety, Laser interlock looks OK 
    Reply  Mon Oct 21 11:44:19 2019, gautam, Update, safety, Interlock reconnected to Innolight controller 
Message ID: 14978     Entry time: Fri Oct 18 18:13:55 2019     Reply to this: 14980
Author: Koji 
Type: Update 
Category: safety 
Subject: Laser interlock looks OK 

I've checked the state of the laser interlock switch and everything looked normal.

ELOG V3.1.3-